40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 6 16:44:37 2011, steve, kiwamu, Update, SUS, BS oplev spectra BS.jpg
    Reply  Wed Jul 6 18:15:26 2011, kiwamu, Update, SUS, Re: BS oplev spectra 
    Reply  Tue Jul 12 17:30:24 2011, steve,, Update, SUS, BS oplev spectra BS_oplev.jpg
       Reply  Thu Jul 14 09:38:50 2011, steve,, Update, SUS, BS oplev spectra BS_oplev.jpg
          Reply  Thu Jul 14 15:27:08 2011, steve,, Update, SUS, SUS oplev spectras 7x
             Reply  Fri Jul 15 09:25:02 2011, rana, Update, SUS, SUS oplev spectras 
Message ID: 4967     Entry time: Thu Jul 14 15:27:08 2011     In reply to: 4966     Reply to this: 4972
Author: steve, 
Type: Update 
Category: SUS 
Subject: SUS oplev spectras 




Healthy BS oplev

 I repeated the BS oplev spectrum today and I do not understand why it does look different. I did it as Kiwamu describes it in entry#4948  The oplev servo was left ON!

 It is working today! Finally I repeated the BS spectra, that we did with Kiwamu last week

 The optical levers were centered during these measurements  without the reference of locked cavities.  They have no reference value now.

SRM sus need some help. ITMX is showing pitch/yaw modes of the pendulum .....OSEM damping is weak?

Attachment 1: BS_oplev.jpg  103 kB  | Hide | Hide all
Attachment 2: PRM_oplev.jpg  131 kB  | Hide | Hide all
Attachment 3: ITMX_oplev.jpg  134 kB  | Hide | Hide all
Attachment 4: ETMX_oplev.jpg  129 kB  | Hide | Hide all
Attachment 5: ETMY_oplev.jpg  102 kB  | Hide | Hide all
Attachment 6: SRM_oplev.jpg  124 kB  | Hide | Hide all
Attachment 7: ITMY_oplev_b.jpg  131 kB  Uploaded Thu Jul 14 16:52:43 2011  | Hide | Hide all
ELOG V3.1.3-